OpenStudy (anonymous):

If a circle has a center at the point (3, 4) and goes through the point (6, 8), what is the radius?

jimthompson5910 (jim_thompson5910):

find the distance between the two points to find the radius

jimthompson5910 (jim_thompson5910):

so you would use the distance formula d = sqrt( (x2-x1)^2 + (y2 - y1)^2 )

OpenStudy (anonymous):

so the answer would be 3

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

-3.32

jimthompson5910 (jim_thompson5910):

d = sqrt( (x2-x1)^2 + (y2 - y1)^2 ) d = sqrt( (6-3)^2 + (8 - 4)^2 ) d = ???